# Diffusion Trajectory-guided Policy for Long-horizon Robot Manipulation

### v1

@@ Abstract
Recently, Vision-Language-Action models (VLA) have advanced robot imitation learning, but high data collection costs and limited demonstrations hinder generalization and current imitation learning methods struggle in out-of-distribution scenarios, especially for long-horizon tasks. A key challenge is how to mitigate compounding errors in imitation learning, which lead to cascading failures over extended trajectories. To address these challenges, we propose the Diffusion Trajectory-guided Policy (DTP) framework, which generates 2D trajectories through a diffusion model to guide policy learning for long-horizon tasks. By leveraging task-relevant trajectories, DTP provides trajectory-level guidance to reduce error accumulation. Our two-stage approach first trains a generative vision-language model to create diffusion-based trajectories, then refines the imitation policy using them. Experiments on the CALVIN benchmark show that DTP outperforms state-of-the-art baselines by in success rate, starting from scratch without external pretraining. Moreover, DTP significantly improves real-world robot performance. All source code and experiment videos will be released upon acceptance.
@@ I Introduction
Imitation Learning (IL) demonstrates significant potential in addressing manipulation tasks within real robotic systems, this is evidenced by its ability to acquire diverse behaviors such as preparing coffee [1] and flipping mugs [2] through learning from expert demonstrations. However, these demonstrations are often limited in coverage, failing to encompass every possible robot pose and environmental variation throughout long-horizon manipulation tasks (Fig. 1(a)). This limitation leads to a key challenge in IL—compounding errors over extended trajectories, where small deviations from the expert trajectory accumulate, ultimately causing task failures. Additionally, robot data is often scarce compared to computer vision tasks because it requires costly and time-consuming human demonstrations. Therefore, improving the generalization capabilities of imitation learning methods using extremely limited and scarce data, given the constraints and high costs of expert demonstrations, becomes a significant challenge.
@@
Recent research has proposed Vision-Language Action (VLA) models [3, 4, 5] to map multi-modality inputs to robot actions by using transformer structures [6]. To guide learning imitation policy, several approaches propose to integrate vision and language to generate a goal image, as seen in methods like Susie [7] or future videos [8, 9], which are pretrained on large-scale video dataset from the Internet. The RT-trajectory [10] uses coarse trajectory sketches as modality instead of language, while the RT-H [11] involves breaking down complex language instructions into simpler, hierarchical commands. For example, instruction as “Close the pistachio jar" can be decomposed step by step into actions like “rotate arm right", “move the arm forward", etc., thereby facilitating robot action generation. These approaches transform complex language instructions into goal images for action generation, replace language commands with coarse trajectory sketches that better align with the action space, or simplify instructions into directional commands that are easier to map to actions. By doing so, they help mitigate compounding errors in imitation policies, particularly in long-horizon tasks. While effective, these methods often rely on manually provided trajectories or goal images, limiting their flexibility, especially in diverse or unstructured environments.
@@
In this paper, we introduce a novel diffusion-based paradigm designed to reduce the feature disparity between the vision-language input and action spaces. By using vision-language input to generate task-relevant 2D trajectories, which are then mapped to the action space, our approach enhances performance in long-horizon robotic manipulation tasks. Unlike robots, which often rely on precise instructions, humans use high-level visualization, such as imagined task-relevant trajectories, to intuitively guide their actions. This visualization aids in adapting to changing conditions and refining our movements in real-time. Similarly, when instructing a robot using language, it should be feasible to envision a task-relevant trajectory to guide the robot’s future actions based on current observations. To facilitate this process, We introduce the Diffusion Trajectory-guided Policy (DTP), which consists of two stages: the Diffusion Trajectory Model (DTM) learning stage and the vision language action policy learning stage. The first stage involves generating a task-relevant trajectory based on a diffusion model. In the second stage, this diffusion trajectory serves as guidance for learning the robot’s manipulation policy, enabling the robot to perform tasks with better data efficiency and improved generalization. We validated our method through extensive experiments on the CALVIN simulation benchmark [12], where it outperformed state-of-the-art baselines by an average success rate of 25% across various settings. Additionally, Our approach is computationally cost-effective requiring only consumer-grade GPUs. The main contributions of the paper include:
@@
We propose the DTP, a novel imitation learning framework that utilizes a diffusion trajectory model to guide policy learning for long-horizon robot manipulation tasks.
@@ II Related Work
Language-conditioned Visual Manipulation Policy Control. Language-conditioned visual manipulation has made significant progress due to advancements in large language models (LLMs) and vision-language models (VLMs). By using task planners like GPT-4 [13] or Palm-E [14], it is possible to break down complex embodied tasks into simpler, naturally articulated instructions. Recently, several innovative methods have been developed in this domain. RT-1 [3] pioneered the end-to-end generation of actions for robotic tasks. RT-2 [4] explores the capabilities of LLMs for Vision-Language-Action (VLA) tasks by leveraging large-scale internet data. RoboFlamingo [15] follows a similar motivation as RT-2, focusing on the utilization of extensive datasets. RT-X prioritizes the accumulation of additional robotic demonstration data to refine training and establish scaling laws in robotic tasks. The Diffusion Policy [2] addresses the prediction of robot actions using a denoising model. Lastly, Octo [16] serves as a framework for integrating the aforementioned contributions into a unified system, further advancing the filed of language-conditioned visual manipulation.
Policy Conditioning Representations. Due to the high-dimensional semantic information contained in language, using video prediction as a pre-training method [9, 17] yields reasonable results. In these approaches, a video prediction model generates future subgoals, which the policy then learns to achieve. Similarly, the goal image generation method [7] utilizes images of subgoals instead of predicting entire video sequences for policy learning. However, both video prediction and goal image generation models often produce hallucinations and unrealistic physical movements. Additionally, these pre-training models demand significant computational resources, posing challenges particularly during inference. RT-trajectory [10] and ATM [18] offer innovative perspectives on generating coarse or particle trajectories, which have proven effective and intuitive. Inspired by these approaches, our method introduces unique adaptations. Unlike RT-trajectory, which generates relatively coarse trajectories through image generation or sketch,and is accompanied by noise with relatively large errors, our method does not completely replace language instructions with coarse trajectories. Instead, we produce high-quality trajectories that can be directly used for end-to-end model inference. Additionally, we use particle trajectories rather than linear trajectories, allowing for more precise and flexible task execution. In contrast to ATM, we model the entire task process using a single key point representing the end-effector’s position in RGB. This key point groundtruth can be readily acquired through the utilization of the camera’s intrinsic and extrinsic parameters. To unify the concept of 2D points or waypoints in the RGB domain, we refer to the sequences of key points from the start to the end of a task as 2D-particle trajectories (Fig. 1(b)). Our method functions similarly to video prediction, serving as a plugin to enhance policy learning.
@@
Diffusion Model for Generation. Diffusion models in robotics are primarily utilized in two areas. Firstly, as previously discussed, they are used for generating future imagery in both video and goal image generation tasks. Secondly, diffusion models are applied to visuomotor policy development, as detailed in recent studies [2, 19, 16]. These applications highlight the versatility of diffusion models in enhancing robotic functionalities. Unlike other methods, our approach does not use diffusion models to directly generate the final policy. Given the high-dimensional semantic richness of language, we propose utilizing diffusion models to create a 2D-particle trajectory. This trajectory represents future end-gripper movements planing in the RGB domain.
@@ III METHOD
Our goal is to create a policy that enables robots to handle long-horizon manipulation tasks by interpreting vision and language inputs. We simplify the VLA task using two distinct phases (Fig. 2(b)(c)): a DTM learning phase and a DTP learning phase. First, we generate diffusion-based 2D particle trajectories for the task. Subsequently, in the second stage, these trajectories are used to guide the learning of the manipulation policy.
@@ III-A Problem Formulation
Multi-Task Visual Robot Manipulation. We consider the problem of learning a language-conditioned policy that take advantage of language instruction , observation , robot states and diffusion trajectory to generate a robot action :
@@ III-B Framework
We introduce the Diffusion Trajectory-guided Policy, as illustrated in Fig. 2. DTP operates within a two-stage framework. In the first stage, our primary focus is on generating the diffusion trajectory which outlines the motion trends essential for completing the task, as observed from a static perspective camera (Fig. 2(b)). This 2D-particle trajectory serves as the guidance for subsequent policy learning. We take a causal transformer as the backbone network which is designed to handle diverse modalities, processing inputs to predict future images and robotic actions with learnable observation and action query tokens respectively. It integrates CLIP [20] as the language encoder for processing language instructions and employs a MAE [21] as the vision encoder for , both of which are with frozen parameters. The vision tokens are then processed with a perceiver resampler [22] to reduce their number. Additionally, it incorporates the robot’s state in world coordinates, as part of its input. All input modalities are shown in Fig. 2(a). Our approach is divided into two main sections. Initially, we detail the process of learning a diffusion trajectory model from the dataset in Section III-C. Subsequently, in Section III-D, we illustrate how diffusion trajectories can be used to guide policy learning for long-horizon robot tasks.
@@ III-C Diffusion Trajectory Model
In the first stage (Fig. 2(b)), we focus on generating diffusion trajectory that maps out the motion trends required for task completion, as viewed from a static perspective camera. To achieve this, we employ a model to transform language instructions and initial visual observations into a sequence of diffusion 2D-particle trajectories . These points indicate the anticipated movements for the remainder of the task:
@@ III-C1 Data Preparation
According to Eq. 2, our input consists of observations and language instruction . For outputs, our aim is to determine the future 2D-particle trajectory of the end effector gripper for finishing the task. Recent advancements in video tracking work make it easy to monitor the end effector gripper [23]. For enhanced convenience and precision, we achieve this by mapping the world coordinates to pixel-level positions according to camera’s intrinsic and extrinsic parameters in the static camera frame, as shown in (Fig. 2(b)) right part. In the first stage, our data format is structured as , facilitating straightforward acquisition of the sequence , thereby simplifying the process of training our model to accurately predict end effector positions.
@@ III-D Diffusion Trajectory-guided Policy
In the second stage, we focus on illustrating how the diffusion trajectory guides the robot manipulation policy (Fig. 2(c)). As previously outlined in our problem formulation, we define our task as a language-conditioned visual robot manipulation task. We base our Diffusion Trajectory-guided Policy on the GR-1 [25] model and incorporate our diffusion trajectory as an additional input, as specified in Eq. 1.
Policy Input. This consists of language and image inputs, as detailed in the Sec. III-B and shown in the left side of Fig. 2(c). To clearly demonstrate our method’s performance, we maintain the same configuration as GR-1. Importantly, for the diffusion trajectory, we do not rely on the inference results from the first training stage. Instead, we use the labeled data from this stage as the diffusion trajectory. This approach enhances precision in training and conserves computational resources by using the labels directly. The simplest training approach is to inject the diffusion particle trajectory directly into the causal baseline. However, our fixed set of 2D particle trajectories can lead to computational intensity during training due to the high number of tokens. Inspired by the perceiver resampler [22], we designed a diffusion trajectory resampler module to reduce the number of trajectory tokens, as shown in Fig. 2(b) and (c).
Diffusion Trajectory as Policy Training. During the policy learning phase (Fig. 2(c)), we generate future particle trajectories to supervise the diffusion trajectory resampler module with . Our policy framework employs a causal transformer architecture, where future particle trajectory tokens are generated prior to action tokens with . This ensures that the particle trajectory tokens effectively guide the formation of action tokens, optimizing the action prediction process in a contextually relevant manner. Additionally, we retain the output of video prediction with , maintaining the same setting as GR-1. This consistency in output makes it easier to conduct ablation studies, as we can directly compare our approach to the original GR-1 model. The optimal DTP objective can be expressed as the following equation:
@@ IV EXPERIMENT
In this section, we evaluate the performance of Diffusion Trajectory Policy on the CALVIN benchmark [12] and real robot.
IV-A CALVIN Benchmark And Baseline
@@
Experiment Setup. we train DTP to predict relative action in positions and Euler angles for arm movements, alongside binary actions for the gripper. The training dataset comprises over 20,000 expert trajectories from four scenes, each paired with language instruction labels. Our DTP method is assessed using the long-horizon benchmark, featuring 1,000 unique sequences of instruction chains articulated in natural language. Each sequence requires the robot to sequentially complete five tasks.
Baselines. We compare our proposed policy against the following state-of-the-art language-conditioned multi-task policies on CALVIN: MT-ACT [26] is a multitask transformer-based policy which predicts action chunk instead of single actions. HULC [27] is a hierarchical approach which predicts latent features of subgoals based on language instructions and observation. RT-1 [3] represents the first approach that utilizes convolutional layers and transformers to generate actions in an end-to-end manner, integrating both language and observational inputs. RoboFlamingo [28] is a fine-tuned Vision-Language Foundation model with 3 billion parameters. It has an additional recurrent policy head specifically designed for action prediction. GR-1 [25] leverages pretraining on the Ego4D dataset, which contains massive-scale human-object interactions captured through web videos. 3D Diffuser Actor [29] integrates 3D scene representations with diffusion objectives to learn robot policies from demonstrations.

### v2

@@ Abstract
Recently, Vision-Language-Action models (VLA) have advanced robot imitation learning, but high data collection costs and limited demonstrations hinder generalization and current imitation learning methods struggle in out-of-distribution scenarios, especially for long-horizon tasks. A key challenge is how to mitigate compounding errors in imitation learning, which lead to cascading failures over extended trajectories. To address these challenges, we propose the Diffusion Trajectory-guided Policy (DTP) framework, which generates 2D trajectories through a diffusion model to guide policy learning for long-horizon tasks. By leveraging task-relevant trajectories, DTP provides trajectory-level guidance to reduce error accumulation. Our two-stage approach first trains a generative vision-language model to create diffusion-based trajectories, then refines the imitation policy using them. Experiments on the CALVIN benchmark show that DTP outperforms state-of-the-art baselines by in success rate, starting from scratch without external pretraining. Moreover, DTP significantly improves real-world robot performance. Our project is at diffusion-trajectory-guided-policy.github.io/.
@@ I Introduction
Imitation Learning (IL) demonstrates significant potential in addressing manipulation tasks within real robotic systems, this is evidenced by its ability to acquire diverse behaviors such as preparing coffee [1] and flipping mugs [2] through learning from expert demonstrations. However, these demonstrations are often limited in coverage [3], failing to encompass every possible robot pose and environmental variation throughout long-horizon manipulation tasks (Fig. 1a)). This limitation leads to a key challenge in IL—compounding errors over extended trajectories, where small deviations from the expert trajectory accumulate, ultimately causing task failures.
Additionally, robot data is often scarce compared to computer vision tasks because it requires costly and time-consuming human demonstrations. Therefore, improving the generalization capabilities of imitation learning methods using extremely limited and scarce data, given the constraints and high costs of expert demonstrations, becomes a significant challenge.
@@
Recent research has introduced Vision-Language Action (VLA) models [4, 5, 6] that map multi-modal inputs to robot actions using Transformer structures [7]. Some approaches, like Susie [8] and others [9, 10], integrate vision and language to generate goal images or future videos, pretrained on large-scale internet datasets. RT-Trajectory [11] uses coarse trajectory sketches instead of language, while RT-H [12] breaks down complex instructions into simpler, hierarchical commands. For instance, "Close the pistachio jar" is decomposed into steps like "rotate arm right" and "move the arm forward," facilitating robot action generation. These methods transform complex instructions into goal images, replace language with trajectory sketches, or simplify instructions into directional commands, mitigating compounding errors in imitation policies, especially in long-horizon tasks. However, they often depend on manually provided trajectories or goal images, limiting flexibility in diverse or unstructured environments.
@@
In this paper, we present a novel diffusion-based paradigm aimed at bridging the feature gap between vision-language inputs and action spaces. By generating task-relevant 2D trajectories from vision-language inputs and mapping them to the action space, our approach improves performance in long-horizon robotic manipulation tasks. Unlike robots that rely on precise instructions, humans use high-level visualizations, like imagined trajectories, to intuitively guide actions, adapting to changes and refining movements in real-time. Similarly, instructing a robot with language should allow envisioning a trajectory to guide future actions based on current observations. To achieve this, we introduce the Diffusion Trajectory-guided Policy (DTP), consisting of two stages: the Diffusion Trajectory Model (DTM) learning stage and the vision-language action policy learning stage, as depicted in Fig. 1c). The first stage generates a task-relevant trajectory using a diffusion model, which then guides the robot’s manipulation policy learning in the second stage, enhancing data efficiency and generalization. The two-stage design allows the first stage to set conditions for the second. We designed the first stage as an independent helper module, easily integrated into any Transformer-based baseline as an additional input, making our approach adaptable across different models. We validated our method through extensive experiments on the CALVIN simulation benchmark [13], achieving a 25% higher average success rate than state-of-the-art baselines across various settings. Additionally, our approach is computationally efficient, requiring only consumer-grade GPUs for training. The main contributions of the paper include:
@@
We propose the Diffusion Trajectory-guided Policy (DTP), a novel imitation learning framework that utilizes a diffusion trajectory model to guide policy learning for long-horizon robot manipulation tasks.
@@ II Related Work
Language-conditioned Visual Manipulation Policy Control. Language-conditioned visual manipulation has made significant progress due to advancements in large language models (LLMs) and vision-language models (VLMs). By using task planners like GPT-4 [14] or PaLM-E [15], it is possible to break down complex embodied tasks into simpler, naturally articulated instructions. Recently, several innovative methods have been developed in this domain. RT-1 [4] pioneered the end-to-end generation of actions for robotic tasks. RT-2 [5] explores the capabilities of LLMs for Vision-Language-Action (VLA) tasks by leveraging large-scale internet data. RoboFlamingo [16] follows a similar motivation as RT-2, focusing on the utilization of extensive datasets. RT-X [17] prioritizes the accumulation of additional robotic demonstration data to refine training and establish scaling laws in robotic tasks. The Diffusion Policy [2] addresses the prediction of robot actions using a denoising model. Lastly, Octo [18] serves as a framework for integrating the aforementioned contributions into a unified system, further advancing the field of language-conditioned visual manipulation.
Policy Conditioning Representations. Leveraging the high-dimensional semantic information in language, video prediction as a pre-training method [10, 19] yields reasonable results by generating future subgoals for the policy to achieve. Similarly, the goal image generation method [8] uses subgoal images instead of full video sequences. However, both approaches often result in hallucinations and unrealistic movements, and they require substantial computational resources, especially during inference. Methods such as those represented by MimicPlay [20] involve learning a latent planner [21]. These approaches require an additional training phase, and the latent planner’s ability to learn useful features can only be indirectly visualized through a decoder, which is not very intuitive. RT-Trajectory [11] and ATM [22] provide innovative methods for generating coarse or particle trajectories. Unlike RT-Trajectory, which uses coarse trajectories with significant noise, we use particle trajectories for generation precision and flexibility. In contrast to ATM, which tracking any sampled points, we use a single key point to illustrate the task process regarding the end-effector’s position in RGB space. This key point is readily derived using the camera’s intrinsic and extrinsic parameters. Motion Tracks [23] and Im2Flow2Act [24] stem from ATM methods. Our method diverges from Motion Tracks by generating language-related long-horizon 2D points, while Motion Tracks only generates short-horizon 2D points without language input. Im2Flow2Act directly uses 2D flow for policy conditions, losing depth information. In contrast, our framework merges the benefits of each method, using 2D points to boost policy output. To standardize the notion of 2D points or waypoints in the RGB space, we label the key point sequences throughout a task as 2D-particle trajectories. Our method functions similarly to video prediction, serving as a plugin to enhance policy learning.
@@
Diffusion Model for Generation. Diffusion models in robotics are primarily utilized in two areas. Firstly, as previously discussed [8, 9, 10], they are used for generating future imagery in both video and goal image generation tasks. Secondly, diffusion models are applied to visuomotor policy development, as detailed in recent studies [2, 25, 18]. These applications highlight the versatility of diffusion models in enhancing robotic functionalities. Unlike these methods, our approach uses diffusion models not to directly generate the final policy but to create a 2D-particle trajectory for future end-gripper movement planning in the RGB domain.
@@ III METHOD
Our goal is to create a policy that enables robots to handle long-horizon manipulation tasks by interpreting vision and language inputs. We simplify the VLA task using two distinct phases (Fig. 2b)c)): a DTM learning phase and a DTP learning phase. First, we generate diffusion-based 2D particle trajectories for the task. Subsequently, in the second stage, these trajectories are used to guide the learning of the manipulation policy.
@@ III-A Problem Formulation
Multi-Task Visual Robot Manipulation. We consider the problem of learning a language-conditioned policy that takes advantage of language instruction , observation , robot states and diffusion trajectory to generate a robot action :
@@ III-B Framework
We introduce the Diffusion Trajectory-guided Policy, as illustrated in Fig. 2. DTP operates within a two-stage framework. In the first stage, our primary focus is on generating the diffusion trajectory which outlines the motion trends essential for completing the task, as observed from a static perspective camera (Fig. 2b)). This 2D-particle trajectory serves as the guidance for subsequent policy learning. We take a causal Transformer as the backbone network which is designed to handle diverse modalities, processing inputs to predict future images and robotic actions with learnable observation and action query tokens respectively. It integrates CLIP [26] as the language encoder for processing language instructions and employs a MAE [27] as the vision encoder for , both of which are with frozen parameters. The vision tokens are then processed with a perceiver resampler [28] to reduce their number. Additionally, it incorporates the robot’s state in world coordinates, as part of its input. All input modalities are shown in Fig. 2a). Our approach is divided into two main sections. Initially, we detail the process of learning a diffusion trajectory model from the dataset in Section III-C. Subsequently, in Section III-D, we illustrate how diffusion trajectories can be used to guide policy learning for long-horizon robot tasks.
@@ III-C Diffusion Trajectory Model
In the first stage (Fig. 2b)), we focus on generating diffusion trajectory that maps out the motion trends required for task completion, as viewed from a static perspective camera. To achieve this, we employ a model to transform language instructions and initial visual observations into a sequence of diffusion 2D-particle trajectories . These points indicate the anticipated movements for the remainder of the task:
@@ III-C1 Data Preparation
According to Eq. 2, our input consists of observations and language instruction . For outputs, our aim is to determine the future 2D-particle trajectory of the end effector gripper for finishing the task. Recent advancements in video tracking work make it easy to monitor the end effector gripper [29]. For enhanced convenience and precision, we achieve this by mapping the world coordinates to pixel-level positions according to camera’s intrinsic and extrinsic parameters in the static camera frame, as shown in (Fig. 2b)) right part. In the first stage, our data format is structured as , facilitating straightforward acquisition of the sequence , thereby simplifying the process of training our model to accurately predict end effector positions.
@@ III-D Diffusion Trajectory-guided Policy
In the second stage, we focus on illustrating how the diffusion trajectory guides the robot manipulation policy (Fig. 2c)). As previously outlined in our problem formulation, we define our task as a language-conditioned visual robot manipulation task. We base our Diffusion Trajectory-guided Policy on the GR-1 [31] model and incorporate our diffusion trajectory as an additional input, as specified in Eq. 1.
Policy Input. This consists of language and image inputs, as detailed in the Sec. III-B and shown in the left side of Fig. 2c). To clearly demonstrate our method’s performance, we maintain the same configuration as GR-1. Importantly, for the diffusion trajectory, we do not rely on the inference results from the first training stage. Instead, we use the labeled data from this stage as the diffusion trajectory. This approach enhances precision in training and conserves computational resources by using the labels directly. The simplest training approach is to inject the diffusion particle trajectory directly into the causal baseline. However, our fixed set of 2D particle trajectories can lead to computational intensity during training due to the high number of tokens. Inspired by the perceiver resampler [28], we designed a diffusion trajectory resampler module to reduce the number of trajectory tokens, as shown in Fig. 2b) and c).
Diffusion Trajectory-guided Policy Training. During the policy learning phase (Fig. 2c)), we generate future particle trajectories to supervise the diffusion trajectory resampler module with . Our policy framework employs a causal Transformer architecture, where future particle trajectory tokens are generated prior to action tokens with . This ensures that the particle trajectory tokens effectively guide the formation of action tokens, optimizing the action prediction process in a contextually relevant manner. Additionally, we retain the output of video prediction with , maintaining the same setting as GR-1. This consistency in output makes it easier to conduct ablation studies, as we can directly compare our approach to the original GR-1 model. The optimal DTP objective can be expressed as the following equation:
@@ IV EXPERIMENT
In this section, we evaluate the performance of Diffusion Trajectory Policy in both simulation and real-world robot experiments.
IV-A CALVIN Benchmark and Baselines
@@
Experiment Setup. We train DTP to predict relative actions in positions and Euler angles for arm movements, along with binary actions for the gripper. Our simulation setup aligns with the base model [31], where both the number of generated and executed actions are set to 1. The training dataset includes over 20,000 expert trajectories from four scenes, each paired with language instruction labels. Our DTP method is evaluated using the long-horizon benchmark, which features 1,000 unique sequences of instruction chains articulated in natural language. Each sequence requires the robot to sequentially complete five tasks.
Baselines. We compare our proposed policy against the following state-of-the-art language-conditioned multi-task policies on CALVIN: MT-ACT [32]: A multitask Transformer-based policy predicting action chunks. HULC [33]: A hierarchical approach predicting latent sub-goal features from language and observations. RT-1 [4]: Utilizes convolutional layers and Transformers for end-to-end action generation from language and observations. RoboFlamingo [34]: A fine-tuned Vision-Language Foundation model with 3 billion parameters. GR-1: Pretrained on the Ego4D dataset, featuring large-scale human-object interactions. 3D Diffuser Actor [35]: Integrates 3D scene representations with diffusion objectives to learn policies from demonstrations.